OpenStudy (anonymous):

What is the sum of this infinite geometric series? 2 + 2/3 + 2/9 + 2/27 + ... a. 2.25 b. 2.5 c. 3 d. infinity e. 10

OpenStudy (anonymous):

I get 3. a = 2, r = 1/3. So a / (1 - r) = 2 / (1 - 1/3) = 3